### v1

# Game Theory in Cosmology

Abstract
We present a game-theoretic statistical framework for cosmology, which we term Cosmological Teleodynamics. We recast the dark sector, cosmic acceleration, large-scale structure, and cosmic tensions as emergent consequences of nonlocal memory and intrinsically persistent organization in a self-gravitating Universe. By introducing a maximum-caliber weight on cosmic histories and a bias functional encoding structural memory, we derive modified Friedmann, Boltzmann, and Poisson equations that naturally generate dark energy-like acceleration, dark matter-like clustering, and scale-dependent growth suppression. We also show how this approach can naturally help alleviate the $H_{0}$ and $S_{8}$ tensions, can produce anisotropic velocity fields, and predict environment-dependent halo signatures that cannot arise from particle dark matter or scalar-field dark energy. We also derive a generalized horizon entropy and temperature, revealing a nonequilibrium statistical origin for cosmic acceleration and formulating a Law of Universal Arbitrage Equilibrium that governs the evolution of the Universe, showing that it is expanding towards a continuous form of Nash equilibrium. Cosmological Teleodynamics therefore offers a unified, emergent, and testable alternative to the conventional dark sector, rooted not in new particles but in the intrinsic statistical and systemic structure of cosmic memory, and shows that the universe itself operates like a giant potential game.
1 Introduction
[…]
Beyond these component mysteries lies a subtler class of phenomena involving gravitational memory and dynamically induced effects. Persistence effects in large-scale structure refer to how cosmic evolution retains information about initial conditions longer than ideal models assume. The web-like structure of the universe — with its filaments, walls, and voids — encodes a remarkably detailed fossil record of primordial density fluctuations and subsequent gravitational dynamics. Gravitational memory effects and dynamically induced preferences emerge when gravitational interactions select particular configurations from initially Boltzmann-weighted distributions. Such alignment effects shape cosmic structure and may constrain modified gravity theories or dark matter properties beyond those accessible through traditional clustering statistics. These subtle phenomena of persistence exemplify how gravity’s cumulative effects give rise to emergent structures that encode rich information about fundamental physics.
To address these concerns, we present a perspective that accounts for persistence effects in a novel manner by using game theory. One might wonder how game theory applies to cosmology. It may seem rather far-fetched. However, as we demonstrate, there is a surprisingly close connection via the concept of persistence. Game theory is concerned with predicting the behavior of goal-driven agents that, by their very nature of incessantly pursuing goals, exhibit persistent behavior. This persistence results in constrained behavior - i.e., persistent agents do not explore the entire phase space but only a part of it. Thus, persistence implies latent constraints on the dynamics and evolution of agents. Game theory offers us a new way to model persistence mathematically as constraints on possible collective behaviors. It is in this sense that we use the word ‘game’ in our theory of cosmology.
In particular, we use a novel game-theoretic framework, statistical teleodynamics, originally developed to model the collective behavior of millions of goal-driven agents in biology, ecology, economics, and sociology [34, 35, 36]. Statistical teleodynamics is a synthesis of the fundamental principles of the theory of potential games [37, 38] and those of statistical mechanics. The name originates from the Greek word telos, meaning ‘goal’. Just as the dynamical behavior of gas molecules in a container is driven by thermal agitation (hence thermodynamics), the dynamics of active agents is driven by their persistent pursuit of goals, and hence teleodynamics. This theory applies to three kinds of goal-driven agents: (i) intentionally persistent (e.g., humans), (ii) instinctively persistent (e.g., bacteria, ants, birds), and (iii) intrinsically persistent (e.g., active colloids, galaxies). In the limiting case of goal-free entities, such as gas molecules in a container, this theory reduces to the familiar results of statistical mechanics.
Thus, we present a teleodynamic model of the universe whose evolution is not purely random but is influenced by persistent structures, memory effects, and dynamically induced preferences. In conventional statistical mechanics, we assume that the probability assigned to a microscopic configuration is determined solely by its energy via the Boltzmann weight. We further assume ergodicity, loss of memory, short-range interactions, etc. These assumptions are questionable for gravitational and cosmological systems where correlations are long-ranged, ergodicity is broken, and macroscopic evolution depends sensitively on prior structure formation. In this sense, teleodynamics expands the statistical toolkit by adding a second structural ingredient, in addition to energy, namely a functional encoding of the system’s persistent organizational bias.
2 Potential Games and Arbitrage Equilibrium
Before we demonstrate the use of game-theoretic concepts in cosmology, we need to take a brief detour to introduce potential games [38]. The theory of potential games focuses on predicting the eventual outcome(s) of a dynamical system consisting of many competing, goal-driven agents, as observed in biology, economics, and sociology. For these games, one can identify a single scalar-valued global function, called a potential ($\phi(\boldsymbol{x})$), which captures the necessary information about the utilities (where $\boldsymbol{x}$ is the state vector of the system). Here, agents continuously pursue arbitrage opportunities to increase their benefits and minimize costs in a competitive environment, thereby optimizing their benefit-cost trade-offs, which we define as effective utility in our theory. The effective utility is essentially a proxy for persistence. Agents persistently seek arbitrage opportunities to increase their effective utility. This collective persistence dynamics results in an equilibrium, known as the Nash equilibrium, when the teleodynamic potential $\phi$ is maximized [38].
The effective utility of an agent, $h_{j}$, in state $j$ is the gradient of potential $\phi(\boldsymbol{x})$, i.e.,
${h}_{j}(\boldsymbol{x})\equiv{\partial\phi(\boldsymbol{x})}/{\partial x_{j}}$ (1)
where $x_{j}=N_{j}/N$ and $\boldsymbol{x}$ is the population vector. $N_{j}$ is the number of agents in the state $j$, and $N$ is the total number of agents. Therefore, we have the following:
$\phi(\boldsymbol{x})$ $=$ $\sum_{j=1}^{n}\int{h}_{j}(\boldsymbol{x}){d}x_{j}$ (2)
where $n$ is the total number of states. To determine the maximum teleodynamic potential, we use the method of Lagrange multipliers with $\mathcal{L}$ as the Lagrangian and $\lambda$ as the Lagrange multiplier for the constraint $\sum_{j=1}^{n}x_{j}=1$:
$\mathcal{L}=\phi+\lambda(1-\sum_{j=1}^{n}x_{j})$ (3)
In equilibrium, all agents enjoy the same utility. That is,
$h_{j}=h^{*}$ (4)
It is an arbitrage equilibrium [39] in which agents no longer have any incentive to switch states as all states provide the same utility $h^{*}$. Thus, the maximization of $\phi$ and $h_{j}=h^{*}$ is equivalent when the equilibrium is unique (i.e., $\phi(\boldsymbol{x})$ is strictly concave [38]).
One can view the thermodynamics of gas molecules in a container via the lens of potential games - the Thermodynamic Game [40]- and define the effective “utility” $h_{j}$ of a molecule in state $j$ as
$h_{j}=-\beta E_{j}-\ln N_{j}$ (5)
where $E_{j}$ is its energy, $N_{j}$ is the number of molecules in state $j$, and $\beta=1/k_{B}T$ is the Boltzmann factor. Using Eq. 5 in Eq. 2, we have
$\phi(\textbf{x})=-\beta\sum_{j=1}^{n}x_{j}E_{j}+\frac{1}{N}\ln\frac{N!}{\prod_{j=1}^{n}(N_{j})!}$ (6)
Using Eq. 6 in Eq. 3, we obtain the well-known Gibbs-Boltzmann exponential probability distribution at equilibrium as
$x_{j}^{*}=\frac{e^{-\beta E_{j}}}{\sum_{j=1}^{n}e^{-\beta E_{j}}}$ (7)
where $x_{j}^{*}=N_{j}^{*}/N$ and $N$ is the total number of molecules (the * denotes the equilibrium outcome). The partition function $Z$ given by
$Z=\sum_{j=1}^{n}e^{-\beta E_{j}}$ (8)
and $x_{j}^{*}$ can also be written as
$x_{j}^{*}=\frac{1}{Z}e^{-\beta E_{j}}$ (9)
Readers familiar with statistical mechanics would immediately recognize that the second term in Eq. 6 corresponds to entropy (SS[35]. Therefore, as observed by Venkatasubramanian [35, 41], we recognize that by maximizing $\phi$, we are indeed maximizing the entropy $S$ subject to the energy constraint given by the first term in Eq. 6.
In general, in statistical teleodynamics, the effective utility of a goal-driven agent is a benefit-cost trade-off function of the form given by [34, 36]
$h_{j}=u_{j}-v_{j}-\ln N_{j}$ (10)
where $u_{j}$ is the benefit, $v_{j}$ is the cost, and $\ln N_{j}$ is the disutility of competition among the agents. Table 1 lists some examples from different domains [41, 36] ($\beta$ represents the Boltzmann factor in thermodynamics, and the other Greek parameters represent different domain-specific concepts for different applications. Similarly, the variables $r_{j},l_{j},H$ and $S_{j}$ represent domain-specific concepts that are not relevant for our present discussion).
The generic teleodynamic potential is given by
$\tilde{\phi}(\boldsymbol{x})=\sum_{j=1}^{n}\int\left[u_{j}-v_{j}-\ln x_{j}\right]dx_{j}$ (11)
The generic teleodynamic partition function is given by
$\tilde{Z}=\sum_{j=1}^{n}\exp(u_{j}-v_{j})$ (12)
and the generic teleodynamic probability distribution is
$x_{j}^{*}=\frac{1}{\tilde{Z}}\exp(u_{j}-v_{j})$ (13)
As shown in Eq. 6, the $\ln N_{j}$ term leads to entropy $S$ upon integration, and the $u_{j}$ and $v_{j}$ terms (upon integration) correspond to the constraints that maximum entropy must satisfy. Thus, the persistence preferences of goal-driven agents, encoded in the terms $u_{j}$ and $v_{j}$, manifest themselves as constraints on maximum entropy. From Table 1, we have (i) for gas molecules, $u_{j}=0$ and $v_{j}=\beta E_{j}$, (ii) for income distribution in economics, $u_{j}=\alpha\ln S_{j}$ and $v_{j}=\eta(\ln S_{j})^{2}$, and (iii) for segregation dynamics in sociology, $u_{j}=\alpha N_{j}+\ln(H-N_{j})$, $v_{j}=\eta N_{j}^{2}$. These different mathematical forms of constraints (i.e., $\sum_{j=1}^{n}\int\left[u_{j}-v_{j}\right]dx_{j}$) for maximum entropy across various domains show how persistence effects can appear mathematically in our framework. Eq. 12 and 13 could get complicated if $u_{j}$ and $v_{j}$ depend on $N_{j}$, as in the case of bird flocking and social segregation games shown in Table 1. However, this is not the case for the Thermodynamic and Income Games. As a result, we can obtain the probability distributions analytically, cleanly - e.g., exponential (energy) and lognormal (income) distributions.
Table 1: Effective utility in different domains
Domain System Effective Utility ($h_{j}$)
Physics Thermodynamics $\beta E_{j}-\ln N_{j}$
Physics Janus Particles $\dfrac{\omega r_{j}^{a}}{a}-\ln N_{j}$
Ecology Ant crater formation $b-\dfrac{\omega r_{j}^{a}}{a}-\ln N_{j}$
Ecology Birds flocking $\alpha N_{j}-\eta N_{j}^{2}+\gamma N_{j}l_{j}-\ln N_{j}$
Sociology Segregation dynamics $\alpha N_{j}-\eta N_{j}^{2}+\ln(H-N_{j})-\ln N_{j}$
Economics Income distribution $\alpha\ln S_{j}-\eta\left(\ln S_{j}\right)^{2}-\ln N_{j}$
3 Statistical Teleodynamics of Cosmology
Guided by the mathematical models of persistence in other domains, as shown in Table 1, we are now ready to develop a teleodynamic framework for cosmology that captures the effects of persistent structures, memory effects, and dynamically induced preferences. We start with Eq. 10, Eq. 12, and Eq. 13, considering a discrete space of microstates $x$. Instead of assigning weights solely by the energy $E(x)$, as in Eq. 5, we introduce an additional functional $\Phi(x)$ which captures a path-dependent or environment-dependent persistence bias as a constraint intrinsic to the system. Therefore, we have $u(x)=0$ and $v(x)=\beta E(x)+\alpha\Phi(x)$, yielding
$h(x)=-\beta E(x)-\alpha\Phi(x)-\ln N(x)$ (14)
The microscopic probability distribution is given by
$p(x)=\frac{1}{Z}\exp[-\beta E(x)-\alpha\Phi(x)]$ (15)
with the partition function
$Z(\beta,\alpha)=\sum_{x}\exp[-\beta E(x)-\alpha\Phi(x)]$ (16)
The parameter $\alpha$ controls the strength of the teleodynamic bias, while the functional $\Phi$ can depend on nonlocal correlations, environmental constraints, dynamical stability conditions, and statistical preferences linked to the history of the system. In the standard thermodynamic limit where $\Phi$ vanishes, one recovers the ordinary canonical ensemble.
Since it is natural to use free energy instead of the game-theoretic (i.e., teleodynamic) potential in cosmology, we will now switch to a free energy-based analysis. As Venkatasubramanian showed [35, 34, 41], minimizing free energy is equivalent to maximizing game-theoretic potential $\phi$.
So, when $\Phi$ is present, the effective free energy becomes
$F_{\text{TD}}=-\frac{1}{\beta}\ln Z=\sum_{x}p(x)E(x)+\frac{1}{\beta}\sum_{x}p(x)\ln p(x)+\alpha\sum_{x}p(x)\Phi(x)$ (17)
This tells us that the equilibrium distribution simultaneously minimizes an energetic, an entropic, and a teleodynamic term. The third contribution has no analog in ordinary thermodynamics and encodes the persistence of organizational constraints and dynamical feedback not reducible to energy alone.
The evolution equations of teleodynamic systems follow by considering the probability flux in the configuration space. So, if $x(t)$ denotes microscopic variables, the coarse-grained dynamics acquires an additional drift proportional to the teleodynamic potential. In a gradient flow approximation, the equation of motion takes the form
$\dot{x}=-\nabla E(x)-\alpha\nabla\Phi(x)+\eta(t)$ (18)
where $\eta$ represents stochastic fluctuations and the deterministic term $\alpha\nabla\Phi$ drives the system toward attractors defined not only by the minima of the energy but also by the minima of the teleodynamic functional. This allows us to naturally generate stable macroscopic tendencies that would require fine-tuning in conventional thermodynamics, and a probability-based formulation gives us a teleodynamic Fokker-Planck equation
$\frac{\partial p(x,t)}{\partial t}=\nabla\cdot\left[p(x,t)\nabla\left(E(x)+\alpha\Phi(x)\right)\right]+D\nabla^{2}p(x,t)$ (19)
demonstrating again that teleodynamics introduces a functional bias, altering both the drift and the stationary solutions.
The structure of this whole theory consists of a few vital ingredients: first is a generalized Gibbs distribution that retains the form of an exponential family but includes an additional functional in the exponent. The second is the appearance of an effective potential $E_{\text{eff}}(x)=E(x)+\alpha\Phi(x)$ that modifies the free energy landscape and the deterministic drift of coarse-grained dynamics. The third is the recognition that complex macroscopic systems may not be described by equilibrium thermodynamics, but rather by a combination of free-energy minimization and attractor selection, governed by $\Phi$. These features become particularly important in cosmology, where long-range gravitational forces, nonlocality, and the presence of large-scale correlated structures imply that the assumptions of ordinary statistical mechanics are not satisfied. In such a context, the teleodynamic corrections provide a natural mechanism for introducing effective energy densities, biasing large-scale flows, and generating attractor behavior without invoking new particle species or additional fundamental fields. The key differences between thermodynamics and teleodynamics are summarized in Table 2.
Property Ordinary thermodynamics Teleodynamics
Microscopic weight $e^{-\beta E(x)}$ $e^{-\beta E(x)-\alpha\Phi(x)}$
Equilibrium criterion minimize free energy minimize $F+\alpha\langle\Phi\rangle$
Assumptions ergodic, memoryless, short range non-ergodic, memory-influenced, nonlocal
Attractors energy minima only minima of $E+\alpha\Phi$
Dynamics $\dot{x}=-\nabla E+\eta$ $\dot{x}=-\nabla(E+\alpha\Phi)+\eta$
Table 2: Key differences between thermodynamics and teleodynamics
4 Cosmological Teleodynamics Framework
The teleodynamic description of cosmological dynamics begins with the recognition that the evolution of a macroscopic system is determined not only by instantaneous energies but also by the statistical weight of its entire history. So we start here by considering the appropriate statistical object to be the maximum–caliber distribution on histories $\Gamma=\{x(t),p(t)\}$ which is defined by
$\mathcal{P}[\Gamma]\propto\exp[-\beta\mathcal{A}[\Gamma]-\alpha\mathcal{K}[\Gamma]]$ (20)
where $\mathcal{A}[\Gamma]$ is the action functional that captures the usual dynamical contributions and $\mathcal{K}[\Gamma]$ is a bias functional that encodes the memory, environmental dependence, and structural correlations that the system retains over time. The parameters $\beta$ and $\alpha$ enforce the macroscopic constraints on energy and teleodynamic bias, which are kind of like Lagrange multipliers of maximum entropy. To make this explicit, we write the mechanical part of the history weight as
$\mathcal{A}[\Gamma]=\int dt\left[\frac{p^{2}}{2a^{2}m}+m\Psi(x,t)\right]$ (21)
where $\Psi$ is the gravitational potential, and the teleodynamic part as
$\mathcal{K}[\Gamma]=\int dt\,\Phi\left(x(t),p(t);f,\mathcal{E}\right)$ (22)
where $\Phi$ is the bias functional depending on the configuration variables, the phase-space distribution $f$, and the coarse environmental fields $\mathcal{E}$ such as the tidal tensor or the surrounding density field. The functional $\Phi$ here accounts for nonlocal correlations, dynamical memory, and persistent environmental structure, and is the core ingredient that distinguishes teleodynamics from ordinary thermodynamics.
The effective equations of motion are derived by extremizing the exponent of $\mathcal{P}[\Gamma]$, and the variational differentiation with respect to $x(t)$ and $p(t)$ yields the drift equations
$\dot{x}=\frac{p}{a^{2}m}$ (23)
$\dot{p}=-\nabla\Psi(x,t)-\alpha\nabla\Phi(x,p,t)-Hp$ (24)
This is helpful as it shows that the dynamics experiences an additional force proportional to the gradient of the bias functional. This force is not sourced from new matter fields but rather represents a coarse-grained structural response of the system to its own past evolution and environment. As a result, the presence of $\alpha\nabla\Phi$ alters the gravitational streaming of matter, modifies the effective potential, and influences the background expansion.
By promoting the trajectory equations to a kinetic description, one obtains the teleodynamic Boltzmann equation, from which the modified continuity, Euler, and growth equations follow. One can decompose the bias functional into its background part and fluctuations,
$\Phi(t,x)=\bar{\Phi}(t)+\varphi(t,x)$ (25)
where $\bar{\Phi}(t)$ contributes a homogeneous energy density to the Friedmann equations and $\varphi(t,x)$ modifies the Poisson equation through an effective clustering density. This decomposition is central, as it provides a view of the dark sector and, therefore, the bias functional and the maximum-caliber weight provide the unifying mathematical structure underlying all teleodynamic effects in cosmology.
4.1 Boltzmann Dynamics and Galactic Phase-space Evolution
The starting point for any analysis of matter evolution in cosmology is the Boltzmann equation of collisionless tracers. The Boltzmann equation considers the evolution of matter for the phase-space distribution $f(t,\mathbf{x},\mathbf{p})$ where $\mathbf{p}=am\mathbf{v}$ denotes the canonical momentum of a collisionless tracer of mass $m$. In comoving coordinates under the assumption that gravity is the only long–range force and that the system possesses no dynamical memory, the Liouville flow allows one to write the standard collisionless Boltzmann equation as [42, 43, 44, 45, 46, 47, 48, 49]
$\frac{\partial f}{\partial t}+\frac{\mathbf{p}}{a^{2}m}\cdot\nabla_{\mathbf{x}}f-\nabla_{\mathbf{x}}\Psi\cdot\nabla_{\mathbf{p}}f-H\,\mathbf{p}\cdot\nabla_{\mathbf{p}}f=0$ (26)
where $\Psi$ is the Newtonian potential determined from the Poisson equation. This form implicitly assumes ergodicity, local relaxation, and the absence of long–lived correlations. However, cosmological structure formation violates all of these assumptions, as matter resides in a cosmic web whose tidal fields and filamentary geometry introduce persistent correlations and memory at all scales. Teleodynamics accommodates this by introducing a nonlocal, memory–carrying drift term. The microscopic bias functional is decomposed as
$\Phi(t,\mathbf{x})=\bar{\Phi}(t)+\varphi(t,\mathbf{x})$ (25)
where $\varphi$ encodes spatially varying correlations and environmental response. The total acceleration then becomes
$\mathbf{a}=-\nabla\Psi-\alpha\,\nabla\Phi$ (27)
so that the teleodynamic Boltzmann equation is
$\frac{\partial f}{\partial t}+\frac{\mathbf{p}}{a^{2}m}\cdot\nabla_{\mathbf{x}}f-\left(\nabla_{\mathbf{x}}\Psi+\alpha\nabla_{\mathbf{x}}\Phi\right)\cdot\nabla_{\mathbf{p}}f-H\,\mathbf{p}\cdot\nabla_{\mathbf{p}}f=0$ (28)
The term proportional to $\alpha$ represents the coarse-grained effect of long-lived correlations, environmental structure, and nonlocal memory encoded by the bias functional. Its presence here reshapes the flow in the phase space and generates an effective potential $\Psi_{\text{eff}}=\Psi+\alpha\Phi$ that governs the trajectories of matter elements. To understand the implications for galactic dynamics, we first compute the velocity moments of this kinetic equation. The continuity equation remains
$\dot{\delta}+\frac{1}{a}\nabla\cdot\left[(1+\delta)u\right]=0$ (29)
but the Euler equation changes to
$\dot{u}+Hu+\frac{1}{a}(u\cdot\nabla)u=-\frac{1}{a}\nabla\Psi-\frac{1}{a(1+\delta)}\nabla\cdot P-\frac{\alpha}{a}\nabla\Phi$ (30)
where $P$ denotes our stress tensor. Linearizing the continuity and Euler equations and transforming to Fourier space leads to the coupled system
$\dot{\delta}(k,t)=-\frac{1}{a}\,\theta(k,t),$ (31)
$\dot{\theta}(k,t)+H\theta(k,t)-4\pi Ga\bar{\rho}\,\delta(k,t)=-\frac{c_{s}^{2}k^{2}}{a}\,\delta(k,t)-\frac{\alpha k^{2}}{a}\,\Phi_{k}(t)$ (32)
showing explicitly that the teleodynamic contribution provides a new scale–dependent source term for the velocity divergence. This term may behave effectively as a renormalization of the gravitational coupling or as a dispersive pressure-like correction depending on the structure of the kernel relating $\Phi_{k}$ to the matter distribution. Depending on the structure of the response $\Phi_{k}$, it is clear that the growth of perturbations acquires either an effective renormalization of the gravitational coupling or an additional pressure-like dispersive term. These corrections propagate directly to the peculiar velocity statistics [50, 51], the velocity-density cross-spectrum, and the bulk flow variance, and in particular, the relation between the divergence of the velocity field and the matter overdensity becomes
$\theta(k,a)=-aHf(k,a)\delta(k,a)$ (33)
This is a scale-dependent growth rate, $f(k,a)$, that departs from the scale independence predicted by the standard collisionless cold dark matter picture.
We see that the teleodynamic modification of the Boltzmann equation, therefore, induces characteristic signatures in galaxy motions and peculiar velocities. The additional drift term produces velocity fields with enhanced coherence over large distances, non-Gaussian velocity distributions, and a partial decorrelation between velocity and density, all of which differ from the predictions of an idealized gravitational gas governed solely by $\Psi$. Observations of bulk flows, pairwise velocities, and redshift-space distortions already hint at such anomalies, suggesting that a framework incorporating nonlocal memory and structural bias may be better suited to describing galactic phase-space dynamics. Unlike the ideal gas approximation, which assumes that particles respond only to instantaneous gravitational forces, the teleodynamic picture accounts for the fact that galaxies evolve within a network of long-range correlations created by the cosmic web and retain dynamical memory of their environment. This makes the teleodynamic Boltzmann framework a more realistic representation of large-scale cosmic motion and establishes the foundation for the remaining ingredients of teleodynamic cosmology.
4.2 Friedmann Dynamics and Dark Energy
The Friedmann equations, as we know, describe the dynamics of an expanding universe in a homogeneous and isotropic FLRW spacetime. In the standard picture, one assumes that the background energy density is given solely by physical matter components so that the expansion obeys
$3M_{P}^{2}H^{2}=\rho,$ (34)
$2M_{P}^{2}\dot{H}=\rho+p$ (35)
Together with the continuity equation
$\dot{\rho}+3H(\rho+p)=0$ (36)
these relations fully determine the FLRW background given a set of matter components with equations of state. One immediate consequence of (34)-(35) is that cosmic acceleration
$\frac{\ddot{a}}{a}>0,$ (37)
[…]
$\Phi(t,x)=\bar{\Phi}(t)+\varphi(t,x)$
with a vanishing spatial mean of the inhomogeneous part and the average $\bar{\Phi}(t)$ enters the background action through an effective contribution to the coarse-grained stress tensor. This leads to a modification of the Friedmann sector to
$3M_{P}^{2}H^{2}=\rho+\alpha\bar{\Phi}(t)$ (38)
and
$2M_{P}^{2}\dot{H}=\rho+p+\alpha\Pi_{\text{TD}}(t)$ (39)
where $\Pi_{\text{TD}}(t)$ signifies the effective pressure associated with the temporal evolution of $\bar{\Phi}$. The structure of these equations shows that the teleodynamic bias contributes an additional homogeneous component to the cosmic energy budget, with the effective density and pressure given by
$\rho_{\text{TD}}(t)=\alpha\bar{\Phi}(t),\quad p_{\text{TD}}(t)=\alpha\Pi_{\text{TD}}(t)$ (40)
Note that no additional fields are required for this term to appear. It follows directly from the statistical and systemic coarse-graining of matter in a gravitationally correlated environment and represents the macroscopic imprint of dynamical memory and nonlocal structure. One can derive (38) and (39) directly by inserting the modified stress tensor
$T^{\mu\nu}_{\mathrm{eff}}=T^{\mu\nu}_{\mathrm{matter}}+T^{\mu\nu}_{\mathrm{TD}}$
with $T^{\mu\nu}_{\mathrm{TD}}$ determined through variational differentiation of the coarse-grained teleodynamic functional
$S_{\mathrm{eff}}[g_{\mu\nu}]=\int d^{4}x\,\sqrt{-g}\left[\mathcal{L}_{\mathrm{matter}}-\alpha\,\Phi\right]$ (41)
Spatial averaging then ensures that only $\bar{\Phi}(t)$ contributes to the background energy density and only $\Pi_{\mathrm{TD}}(t)$ contributes to the effective pressure.
The condition for accelerated expansion is obtained by combining the Friedmann equations in the usual way, as
$\frac{\ddot{a}}{a}=-\frac{1}{6M_{P}^{2}}[\rho+3p+\rho_{\text{TD}}+3p_{\text{TD}}]>0$ (42)
In the absence of teleodynamic corrections, this inequality demands an exotic fluid with $p\simeq-\rho$, which is the basis of the dark energy. But with teleodynamic contributions accounted for correctly, the accelerating condition may be satisfied even if the physical matter components obey normal equations of state. Furthermore, when the bias functional is sufficiently slowly varying, the effective equation of state satisfies
$w_{\text{TD}}\equiv\frac{p_{\text{TD}}}{\rho_{\text{TD}}}=\frac{\Pi_{\text{TD}}(t)}{\bar{\Phi}(t)}\approx-1$ (43)
and the teleodynamic term behaves like a cosmological constant. More generally, the rate of change of $\bar{\Phi}$ can also determine whether the expansion is quintessence, phantom, or quintom-like. This shows that the entire late-time acceleration can be interpreted simply as a large-scale statistical and systemic effect of a memory-bearing bias functional in a self-gravitating universe.
Instead of introducing new scalar fields, potentials, or hidden matter sectors in our theory [52, 53, 54], one simply recognizes that the universe is not a system that relaxes to a gas-like equilibrium or one whose expansion is governed solely by local microphysics. The long-range nature of gravity and the persistent structures generated by cosmic evolution imply that the background expansion should include contributions that have nonlocal correlations and path dependence. The teleodynamic energy density $\rho_{\text{TD}}$ arises precisely from these considerations and provides an emergent explanation for the cosmic acceleration. The term $\rho_{\mathrm{TD}}=\alpha\bar{\Phi}$ arises not from a new microscopic field but from the coarse-grained effect of self-gravitating correlations and memory, which accumulate as the Universe forms persistent structures and departs more strongly from ergodicity. In this sense, we see that late-time acceleration emerges naturally as the Universe becomes increasingly non-thermal and the expansion is governed by an effective potential generated by the ensemble of long-lived gravitational correlations. Teleodynamics in this sense provides a statistical and systemic explanation for cosmic acceleration grounded in the physics of memory-bearing, nonlocally correlated matter rather than in additional dark sectors or fine-tuned vacuum energies.
4.3 Dark Matter and Structure Growth
In standard cosmology, the problem of dark matter is addressed by introducing a new, non-relativistic matter component whose density obeys the usual continuity and Euler equations and whose gravitational potential satisfies the Poisson equation. Teleodynamics now offers a different perspective in which the missing mass effect arises not from a new particle species or primordial remnants, but from the structural correction induced by the teleodynamic bias functional at the level of inhomogeneities. This emerges naturally once one decomposes the bias into its homogeneous and fluctuating parts again, as before
$\Phi(t,x)=\bar{\Phi}(t)+\varphi(t,x)$
and allows the inhomogeneous component to participate in the gravitational potential felt by matter elements.
The teleodynamic modification of the kinetic theory would imply that particles experience an effective potential consisting of the Newtonian part plus a correction generated by the bias functional. This potential is defined by
$\Psi_{\text{eff}}(t,x)=\Psi(t,x)+\alpha\varphi(t,x)$ (44)
with the gravitational potential obeying the usual Poisson equation
$\nabla^{2}\Psi=4\pi Ga^{2}\bar{\rho}\delta$ (45)
To determine the additional contribution, one applies the Laplacian to $\Psi_{\text{eff}}$, giving us
$\nabla^{2}\Psi_{\text{eff}}=4\pi Ga^{2}\bar{\rho}\delta+\alpha\nabla^{2}\varphi$ (46)
This motivates the definition of an emergent clustering density, which would be
$\rho_{\text{TD}}(t,x)=\frac{\alpha}{4\pi Ga^{2}}\nabla^{2}\varphi(t,x)$ (47)
so that the effective potential satisfies the Poisson equation
$\nabla^{2}\Psi_{\text{eff}}=4\pi Ga^{2}[\bar{\rho}\delta+\rho_{\text{TD}}]$ (48)
Interestingly, the term $\rho_{\text{TD}}$ acts as a source of gravitational clustering, playing the role conventionally attributed to dark matter. Its origin is again statistical and systemic, rather than material. It is derived from the curvature of the teleodynamic bias, which captures the influence of long-range correlations, tidal structure, or flow-dependent memory on the local gravitational environment. No new fields or particles are required for this term to emerge, and so it is an effective density derived from the nonlocal structure of the cosmic matter distribution.
The dynamical influence of $\rho_{\text{TD}}$ is made more explicit by examining the Euler equation for the peculiar velocity field derived from the teleodynamic Boltzmann equation, in which, after linearization in Fourier space, we can find that
$\dot{\theta}(k,t)+H\theta(k,t)-4\pi Ga\bar{\rho}\delta(k,t)=-\frac{\alpha k^{2}}{a}\varphi_{k}(t)$ (49)
which, upon using the definition of the effective density, becomes
$\dot{\theta}(k,t)+H\theta(k,t)-4\pi Ga[\bar{\rho}\delta(k,t)+\rho_{\text{TD}}(k,t)]=0$ (50)
This shows that the teleodynamic term enhances the effective gravitational strength precisely as required to explain galactic rotation curves, cluster dynamics, and gravitational lensing. The enhancement is determined by the response kernel that relates $\varphi_{k}$ to matter fluctuations and their tidal environment, so the scale and time dependence of the effective clustering can be adapted to the observed phenomenology without invoking any particulate dark matter. The phenomenological versatility here becomes apparent once one specifies the form of the response kernel connecting $\varphi_{k}$ with the underlying matter perturbations, as in general we can write $\varphi_{k}=K(k,a)\delta_{k}$, with $K(k,a)$ signifying nonlocal environmental memory, tidal couplings, or velocity–space correlations. The effective density then takes the form $\rho_{\text{TD}}(k,a)=\Delta\mu(k,a)\bar{\rho}\delta(k,a)$, where $\Delta\mu(k,a)\equiv\alpha K(k,a)/(4\pi Ga^{2})$. The scale and time dependence of $K$ determines the strength of clustering and thereby allows teleodynamic dark matter to mimic a wide range of popular dark matter phenomenologies.
For example, a kernel that tends to a constant on large scales but grows mildly toward smaller scales reproduces the behavior of cold dark matter, in which the gravitational coupling is nearly scale-independent across the observed linear regime. A kernel with a transition scale $k_{\text{fs}}\sim\ell_{\text{env}}^{-1}$, reflecting the characteristic environmental length $\ell_{\text{env}}$, generates suppressed clustering for $k>k_{\text{fs}}$ and thus imitates warm dark matter without invoking relics. If the kernel decreases for wavenumbers associated with dense environments, then $K$ induces a pressure-like contribution to the Euler equation, potentially yielding cored halo profiles reminiscent of self-interacting dark matter models. A kernel with an oscillatory or rapidly varying structure at small $k$ produces an effective scale-dependent modification to the force law that generates ultralight scalar dark matter effects.
The teleodynamic interpretation of dark matter is hence closely linked to the non-ergodic and nonlocal character of gravitational dynamics. In a universe where structure formation generates filaments, walls, and tidal fields extending over tens or hundreds of megaparsecs, the local gravitational force experienced by a galaxy cannot depend solely on the instantaneous Newtonian potential sourced by visible matter. So, the memory of the surrounding cosmic web, which is encoded in the bias functional, will contribute an additional drift proportional to $\nabla\Phi$, which manifests itself as an effective mass distribution. The emergent density $\rho_{\text{TD}}$ will reflect this coarse-grained structural correction and provide a natural explanation for the observed gravitational anomalies. Unlike standard dark matter, the teleodynamic approach attributes the missing mass influence to the statistical system structure of the cosmic environment, and the gravitational field responds not only to the local matter density but also to the nonlocal bias generated by correlations in the cosmic web.
What is appealing here is that with teleodynamics, we have a unified interpretation on hand. This interpretation of dark matter follows from the same teleodynamic principles that reproduce dark energy at the background level, illustrating that the dark sector can emerge from the universe’s intrinsic statistical and systemic behavior rather than from hidden particle fields. In this way, the teleodynamic framework provides a conceptually “economical” and dynamically consistent account theory for both dark matter and dark energy.
4.4 Explanations of $H_{0}$ and $S_{8}$ tensions
The persistent discrepancy between early and late-time measurements of the Hubble constant $H_{0}$ together with the suppressed amplitude of matter clustering quantified by $S_{8}$, suggests that the standard cosmological model may be missing an element that influences both the background expansion and the growth of structure. Teleodynamics naturally provides such an element, since the background value of the bias functional modifies the Friedmann equations, while its inhomogeneous component alters the gravitational clustering on sub-horizon scales, as we have discussed so far. The background modification stems from the teleodynamic energy density
$\rho_{\text{TD}}(t)=\alpha\bar{\Phi}(t)$ (51)
which enters the Friedmann equation through
$3M_{P}^{2}H^{2}=\rho_{m}+\rho_{r}+\rho_{\text{TD}}$ (52)
If $\bar{\Phi}(t)$ evolves slowly at late times, then $\rho_{\text{TD}}$ behaves like a DE-like component, as we had discussed
$w_{\text{TD}}(t)=\frac{\alpha\Pi_{\text{TD}}(t)}{\alpha\bar{\Phi}(t)}\simeq-1+\delta w(t)$
A small deviation $\delta w(t)$ from 1-1 at late redshifts modifies the distance–redshift relation in a manner that can shift the inferred value of $H_{0}$ upward without altering the physics of recombination. Such a modification is precisely what is required to reconcile the local distance ladder measurements with the CMB inferred value. Writing the Friedmann equation in the form
$H^{2}(z)=H_{0}^{2}\left[\Omega_{m}(1+z)^{3}+\Omega_{r}(1+z)^{4}+\Omega_{\text{TD}}(z)\right]$ (53)
with
$\Omega_{\text{TD}}(z)=\frac{\rho_{\text{TD}}(z)}{3M_{P}^{2}H_{0}^{2}}$ (54)
one sees that the effective DE density inherits the redshift dependence of $\bar{\Phi}(t)$ and if $\bar{\Phi}$ grows mildly at late times, then $\Omega_{\text{TD}}(z)$ exceeds a constant $\Lambda$ behavior below $z\lesssim 0.5$. The result is a slightly faster late universe expansion rate, which reduces the predicted luminosity distance
$D_{L}(z)=(1+z)\int_{0}^{z}\frac{dz^{\prime}}{H(z^{\prime})}$ (55)
at fixed $H_{0}$. When one fits this modified distance-redshift relation to supernova data, the preferred value of $H_{0}$ should shift upward to compensate for the reduced distance integral, and the key point is that this shift arises without modifying $H(z)$ at recombination and therefore leaves the CMB inferred value of the sound horizon untouched.
This behavior becomes clearer when one expands the teleodynamic energy density around the present epoch, as we can write
$\rho_{\text{TD}}(z)=\rho_{\text{TD},0}\left[1+\epsilon_{1}z+\epsilon_{2}z^{2}+\cdots\right]$ (56)
with $\epsilon_{1},\epsilon_{2}$ determined by the slow evolution of $\bar{\Phi}$ and the Hubble parameter near $z=0$ becomes
$H^{2}(z)=H_{0}^{2}\left[1+\frac{3}{2}\Omega_{m}z+\Omega_{\text{TD},0}(\epsilon_{1}z+\epsilon_{2}z^{2}+\cdots)\right]$ (57)
A positive $\epsilon_{1}$ produces an enhancement in $H(z)$ for $0<z\lesssim 0.2$, which is precisely the redshift regime probed by local supernovae. This gives an observational degeneracy between $(H_{0},\epsilon_{1})$ as an increase in $\epsilon_{1}$ shifts the preferred $H_{0}$ upward until the predicted $D_{L}(z)$ matches the supernova Hubble diagram. Because $\epsilon_{1}$ is controlled by the statistical evolution of $\bar{\Phi}$, the required shift in $H_{0}$ arises naturally from teleodynamic considerations. As a result, the $H_{0}$ tension emerges as a natural indication of the non-equilibrium and memory-bearing character of cosmic expansion rather than a failure of the standard matter-radiation sector.
The second component of the teleodynamic correction affects the growth of structure, as the effective potential experienced by matter perturbations is, as discussed before, to be
$\Psi_{\text{eff}}=\Psi+\alpha\varphi$
and its Poisson equation reads
$\nabla^{2}\Psi_{\text{eff}}=4\pi Ga^{2}[\bar{\rho}\delta+\rho_{\text{TD}}]$
Writing the emergent clustering density in Fourier space as
$\rho_{\text{TD}}(k,a)=\Delta\mu(k,a)\bar{\rho}\delta(k,a)$
the linear growth equation becomes
$\ddot{\delta}(k,a)+2H\dot{\delta}(k,a)-4\pi G\bar{\rho}[1+\Delta\mu(k,a)]\delta(k,a)=0$ (58)
A negative or scale-dependent contribution $\Delta\mu(k,a)$ would suppress the gravitational coupling on the weak lensing and redshift space distortion scales, which would end up reducing the amplitude of clustering. This naturally leads to a lower value of $S_{8}$, which is defined through
$S_{8}=\sigma_{8}\sqrt{\frac{\Omega_{m}}{0.3}}$ (59)
This is done without requiring modifications to early universe physics or introducing new particle species. The suppression can arise from the structure of the response kernel $\varphi_{k}$, which depends on the tidal environment or the history of matter flows. In this way, the teleodynamic correction provides the desired reduction in the growth rate, while leaving early-time observables unchanged.
The crucial point is that the same teleodynamic bias functional $\Phi(t,x)$ generates both of these corrections. Its homogeneous part $\bar{\Phi}(t)$ adjusts the late-time expansion rate and yields an effective dark energy component capable of raising the inferred value of $H_{0}$, while its inhomogeneous part $\varphi(t,x)$ modifies the gravitational potential at the level of perturbations, reducing the growth of the structure and producing a lower $S_{8}$. We again see a valuable result: no new exotic particles, additional fields, or modifications of early universe cosmology are required. Both phenomena emerge from the statistical and systemic behavior of a non-ergodic, gravitationally correlated cosmic system.
To make it clear how the teleodynamic approach can address the tensions, we now explicitly show that even a single teleodynamic bias functional can simultaneously shift the inferred value of $H_{0}$ and suppress the late–time growth of structure, thus lowering $S_{8}$. The key point to note again is that the bias functional naturally separates into a homogeneous and an inhomogeneous part, $\Phi(z,\mathbf{x})=\bar{\Phi}(z)+\varphi(z,\mathbf{x})$ and both parts grow as $z\to 0$ in accordance with the teleodynamic principle that memory accumulates as the Universe evolves and structure becomes increasingly persistent. We shall therefore consider a functional form for $\Phi$ constructed in inverse powers of $(1+z)$ so that memory remains negligible at early times, but builds coherently at late times. We begin with the homogeneous part where we can write the teleodynamic energy density as $\rho_{\rm TD}(z)=\alpha\,\bar{\Phi}(z)$ and a natural late–time memory accumulating functional is
$\bar{\Phi}(z)=\bar{\Phi}_{0}\left[1+\epsilon\left((1+z)^{-n}-1\right)\right]$ (60)
where $\epsilon>0$ and $n>0$ quantify the growth of teleodynamic memory. This form satisfies the physical requirement that memory vanishes for $z\gg 1$ and increases monotonically as $z\to 0$. The corresponding teleodynamic density becomes
$\rho_{\rm TD}(z)=\rho_{\rm TD,0}\left[1+\epsilon\left((1+z)^{-n}-1\right)\right]$ (61)
where $\rho_{\rm TD,0}=\alpha\bar{\Phi}_{0}$. If we now define $\Omega_{\rm TD,0}=\rho_{\rm TD,0}/(3M_{P}^{2}H_{0}^{2})$, the Friedmann equation may be written as
$\frac{H^{2}(z)}{H_{0}^{2}}=\Omega_{m}(1+z)^{3}+\Omega_{r}(1+z)^{4}+\Omega_{\rm TD,0}\left[1+\epsilon\left((1+z)^{-n}-1\right)\right]$ (62)
To extract the low redshift behaviour relevant for distance ladder determinations of $H_{0}$, we expand $(1+z)^{-n}$ for small $z$
$(1+z)^{-n}\simeq 1-nz+\frac{n(n+1)}{2}z^{2}+\cdots$ (63)
Inserting this into the Friedmann equation and using spatial flatness, $\Omega_{m}+\Omega_{r}+\Omega_{\rm TD,0}=1$, gives
$\frac{H^{2}(z)}{H_{0}^{2}}\simeq 1+\left(3\Omega_{m}-\epsilon n\Omega_{\rm TD,0}\right)z+\mathcal{O}(z^{2})$ (64)
[…]
$H(z)\simeq H_{0}\left[1+\frac{1}{2}\left(3\Omega_{m}-\epsilon n\Omega_{\rm TD,0}\right)z\right]$ (65)
one obtains
$\frac{1}{H(z)}\simeq\frac{1}{H_{0}}\left[1-\frac{1}{2}\left(3\Omega_{m}-\epsilon n\Omega_{\rm TD,0}\right)z\right]$ (66)
[…]
$H_{0}^{\rm fit}\simeq H_{0}^{\rm true}\left[1+\frac{1}{4}\epsilon\,n\,\Omega_{\rm TD,0}\langle z\rangle\right]$ (67)
where $\langle z\rangle$ denotes a characteristic supernova redshift. A moderate choice such as $\epsilon\sim 2$ and $n\sim 1$ is sufficient to generate an upward shift of $6$-$8\%$ in the inferred $H_{0}$, thus easing the Hubble tension without modifying conditions at recombination.
We now consider the inhomogeneous part of the bias functional and its influence on the growth of the structure. To model the accumulation of environmental memory at late times, we adopt the functional form
$\Delta\mu(k,z)=-\mu_{0}\,(1+z)^{-p}\exp\left[-\left(\frac{k}{k_{*}}\right)^{2}\right]$ (68)
with $\mu_{0}>0$, $p>0$, and characteristic scale $k_{*}$. The factor $(1+z)^{-p}$ ensures that the inhomogeneous teleodynamic memory is negligible at early times but increases as $z\to 0$, while the Gaussian selects the scales at which the persistent tidal structure is most relevant for growth suppression, and so the linear growth equation then becomes
$\ddot{\delta}(k,z)+2H\dot{\delta}(k,z)-4\pi G\,\bar{\rho}(z)\left[1+\Delta\mu(k,z)\right]\delta(k,z)=0$ (69)
For modes near the scale $k\sim k_{*}$ and small $\Delta\mu$, approximate power law solutions $\delta\propto a^{s}$ can satisfy
$s\simeq 1+\frac{3}{5}\Delta\mu_{\rm eff}$ (70)
[…]
$\ln\left(\frac{D_{\rm TD}}{D_{\Lambda{\rm CDM}}}\right)\simeq\frac{3}{5}\Delta\mu_{\rm avg}\ln 2$ (71)
where $\Delta\mu_{\rm avg}$ is a time-averaged value of $\Delta\mu$ over $0<z<1$. A suppression of order 10% in the growth factor, which would be sufficient to reduce $S_{8}$ from $\sim 0.83$ to $\sim 0.75$ requires $\Delta\mu_{\rm avg}\sim-0.2$. Choosing, for example, $\mu_{0}\sim 0.3$, $p\sim 2$, and $k_{*}\sim 0.15\,h\,{\rm Mpc}^{-1}$ produces
$\Delta\mu(z)\in[-0.3,-0.075]$ (72)
over this redshift range, giving us the required average. We can finally now combine the homogeneous and inhomogeneous components; the full bias functional takes the form
$\Phi(z,\mathbf{x})=\bar{\Phi}_{0}\left[1+\epsilon\left((1+z)^{-n}-1\right)\right]+\mu_{0}(1+z)^{-p}\int d^{3}k\,e^{-(k/k_{*})^{2}}\,\delta(k,z)\,e^{i\mathbf{k}\cdot\mathbf{x}}$ (73)
The first term increases slowly as $z\to 0$ and induces a late time uplift of $H(z)$ that raises the inferred $H_{0}$, while the second term strengthens at low redshift and introduces a scale-dependent reduction in the effective gravitational coupling that suppresses the growth of structure on the scales relevant to $S_{8}$. In this way, we see that both tensions arise as complementary manifestations of the same teleodynamic memory accumulation mechanism, and that no new matter fields or modifications of early-time physics are required.
4.5 Teleodynamic Cosmological Entropy
The concept of cosmic horizon entropy in standard cosmology is derived from the Bekenstein-Hawking law [55, 56, 57, 58, 59, 60, 61, 62, 63, 64, 65, 66, 67, 68, 69, 70, 71]. Teleodynamics provides a different route in which the entropy associated with the cosmological horizon can emerge from a statistical counting of coarse-grained microstates on the apparent horizon, without assuming thermodynamic equilibrium or invoking a pre-assigned temperature.
To construct the entropy from teleodynamic first principles, we can model the horizon as a 2-dimensional spherical surface of radius $r_{H}=H^{-1}$ with area $A=4\pi H^{-2}$. The microscopic degrees of freedom associated with this surface are coarse-grained into independent cells of characteristic correlation length $\ell_{\chi}$. The number of effectively independent cells is therefore given as
$N(H)=\frac{A}{\ell_{\chi}^{2}}=\frac{4\pi}{\ell_{\chi}^{2}H^{2}}$ (74)
The state space of each cell is governed by a teleodynamic probability distribution of the form given by
$p_{i}=\frac{1}{z_{\text{cell}}}\exp[-\lambda_{\tiny E}E_{i}-\lambda_{\tiny\Phi}\Phi_{i}]$ (75)
where $E_{i}$ is an effective surface energy of the cell and $\Phi_{i}$ is the local value of the teleodynamic bias functional. So, we have the horizon partition function, then factorizing as
$Z(H)=[z_{\text{cell}}(\lambda_{E},\lambda_{\Phi})]^{N(H)}$ (76)
The entropy of the full horizon is obtained from the Shannon-Gibbs expression
$S_{\text{\tiny TD}}(H)=-\sum_{\{i\}}p_{i}\ln p_{i}=N(H)s_{\text{cell}}(\lambda_{\tiny E},\lambda_{\tiny\Phi})$ (77)
Note here that $s_{\text{cell}}$ denotes the entropy per cell and inserting the expression for $N(H)$ gives us
$S_{\text{\tiny TD}}(H)=\frac{4\pi s_{\text{cell}}}{\ell_{\chi}^{2}}\frac{1}{H^{2}}\equiv\frac{C_{\text{\tiny TD}}}{H^{2}}$ (78)
with the teleodynamic entropy coefficient
$C_{\text{\tiny TD}}=\frac{4\pi s_{\text{cell}}}{\ell_{\chi}^{2}}$ (79)
This is the fundamental area-scaling law of teleodynamic cosmological entropy. It resembles the Bekenstein-Hawking form in that $S\propto A$, but its normalization is not fixed by quantum gravity input, but instead depends on the coarse-graining scale $\ell_{\chi}$ and on the teleodynamic parameters encoded in $s_{\text{cell}}$.
Since the Universe is not in true thermodynamic equilibrium and $H(t)$ evolves in time, the entropy also receives a non-equilibrium correction arising from entropy production associated with the teleodynamic drift. So in the maximum-caliber formalism, the rate of entropy production takes the general quadratic form
$\sigma_{\text{\tiny TD}}(t)=\sum_{a}L_{ab}(t)J_{a}(t)J_{b}(t)\geq 0$ (80)
where $J_{a}$ are coarse-grained fluxes of the horizon microstates driven by the teleodynamic functional, and $L_{ab}$ are generalized Onsager coefficients. If we now integrate this non-equilibrium production term, it gives us the full teleodynamic horizon entropy
$S_{\text{\tiny TD}}(t)=\frac{C_{\text{\tiny TD}}}{H(t)^{2}}+\int^{t}\sigma_{\text{\tiny TD}}(t^{\prime})dt^{\prime}$ (81)
The first term is the geometric area contribution, while the second is the history-dependent memory contribution. This decomposition distinguishes teleodynamic entropy from the purely geometric Bekenstein-Hawking form, revealing the influence of nonlocal and non-ergodic microphysics that is absent from an equilibrium derivation. We can also get the temperature associated with the horizon in this framework, which emerges not from quantum field theory but from the statistical definition
$T^{-1}(t)=\frac{\partial S_{\text{\tiny TD}}}{\partial E_{H}}$ (82)
where $E_{H}$ is the total energy inside the horizon. In standard cosmology, this energy is
$E_{H}(t)=\frac{4\pi}{3}r_{H}^{3}\rho_{\text{eff}}=\frac{4\pi M_{P}^{2}}{H(t)}$ (83)
after using the Friedmann equation $\rho_{\text{eff}}=3M_{P}^{2}H^{2}$ and differentiating with respect to $H$ gives us
$\frac{dE_{H}}{dH}=-\frac{4\pi M_{P}^{2}}{H^{2}},\quad\frac{dS_{\text{\tiny TD}}}{dH}=-\frac{2C_{\text{\tiny TD}}}{H^{3}}$ (84)
Neglecting the small derivative of the history term during slowly varying stages gives the emergent temperature as
$T_{\text{\tiny TD}}(H)=\left(\frac{\partial S_{\text{\tiny TD}}}{\partial E_{H}}\right)^{-1}=\frac{2\pi M_{P}^{2}}{C_{\text{\tiny TD}}}H$ (85)
Thus the temperature scales linearly with $H$, but its normalization differs from the Gibbons–Hawking value unless $C_{\text{TD}}$ takes the special value
$C_{\text{\tiny TD}}=8\pi^{2}M_{P}^{2}$ (86)
[…]
4.6 Observational Signatures of Cosmological Teleodynamics
We next consider the observational consequences of teleodynamics. Since this framework is based on nonlocal memory and environmental correlations rather than additional particle fields, several of its signatures cannot be reproduced by conventional dark sector models, which is an advantage. These include distinctive patterns in galactic halo profiles, anomalous behavior of peculiar velocities and scale-dependent modifications in large scale structure growth.
To start, galactic halos display several features that are difficult to reconcile with conventional particle dark matter models, including the regularity of rotation curves, the diversity of halo shapes, the stability of cores in low-mass systems, and the apparent environmental dependence of halo concentration. In the teleodynamic perspective, the effective potential that governs the halo dynamics is
$\Psi_{\text{eff}}=\Psi+\alpha\varphi$
The additional term reflects the influence of the cosmic web on galactic motions. Because $\varphi$ arises from large-scale tidal fields and nonlocal memory, the resulting gravitational field acquires a component that aligns with filaments and walls, which can lead to producing anisotropic forces that any cold or warm dark matter particle cannot generate. This leads to observable signatures, such as systematically aligned halo axes, environmentally modulated rotation curve slopes, and departures from universal density profiles [72, 73, 74, 75, 76, 77]. Early indications of such effects appear in the observed connection between galactic rotation curves and their large-scale environment, as well as in the increasing evidence for anisotropic motions within galaxy groups [78, 79, 80, 81, 82, 83, 84, 85]. The anisotropic character of the teleodynamic correction becomes explicit once one decomposes the bias fluctuation into scalar, vector and tensor tidal components, as
$\varphi(x)=T^{ij}(x)\frac{\partial_{i}\partial_{j}}{\nabla^{2}}\delta(x)$ (87)
with $T^{ij}$ denoting the large-scale tidal tensor aligned with the cosmic web. The teleodynamic force contribution becomes
$\mathbf{F}_{\text{\tiny TD}}(x)=-\alpha\nabla\varphi(x)=-\alpha T^{ij}\nabla_{i}\left[\frac{\partial_{j}}{\nabla^{2}}\delta(x)\right]$ (88)
Since $T^{ij}$ is anisotropic and inherits the filamentary geometry of the distribution of the surrounding matter, the resulting acceleration is direction-dependent. It is to be noted that no collisionless dark matter particle species can generate a force term of this form, since CDM and WDM respond only to the scalar Newtonian potential. The effective potential can also be written as
$\Psi_{\text{eff}}(x)=\Psi(x)+\alpha T^{ij}\frac{\partial_{i}\partial_{j}}{\nabla^{2}}\delta(x)$ (89)
so that the elliptical nature of halo potentials satisfies
$\nabla^{2}\Psi_{\text{eff}}\propto\delta+\alpha T^{ij}\hat{k}_{i}\hat{k}_{j}\delta(k)$ (90)
which predicts aligned halo axes and anisotropic mass distributions correlated with the cosmic web orientation. Such a directional dependence is absent in standard dark matter models, but it arises naturally in teleodynamics because of the bias functional.
Peculiar velocities offer a second window into teleodynamic effects, as the relation
$\theta(k,a)=-aHf(k,a)\delta(k,a)$ (91)
is modified when the gravitational potential contains the term $\alpha\varphi$, as this modification alters the force law rather than the matter content, it generates a scale-dependent growth rate
$f(k,a)=\frac{d\ln D(k,a)}{d\ln a}$ (92)
[…]
$\theta(k,a)=-aH[1+\Delta\mu(k,a)]\delta(k,a)$ (93)
where $\Delta\mu(k,a)=\alpha K(k,a)/(4\pi G)$ encodes the teleodynamic enhancement or suppression of the effective gravitational coupling, and the presence of $\Delta\mu(k,a)$ introduces a directional dependence in the velocity field whenever the kernel couples to tidal tensors or coherent large-scale structures. In real space, this generates a velocity field of the form
$u_{i}(x,a)=-aH\int\frac{d^{3}k}{(2\pi)^{3}}e^{ik\cdot x}\left[\hat{k}_{i}+\Delta\mu(k,a)T_{ij}(\hat{k})\hat{k}_{j}\right]\frac{\delta(k,a)}{k}$ (94)
where $T_{ij}$ is the local tidal tensor as well. This expression shows that peculiar velocities here receive anisotropic contributions aligned with filaments and sheets of the cosmic web, which is an effect impossible to reproduce with any isotropic particle dark matter model.
These anisotropic corrections propagate directly into the bulk-flow statistics, and the variance of the smoothed velocity field becomes
$\langle u^{2}(R)\rangle=a^{2}H^{2}\int\frac{d^{3}k}{(2\pi)^{3}}[1+\Delta\mu(k,a)]^{2}\frac{P_{\delta\delta}(k,a)}{k^{2}}W^{2}(kR)$ (95)
[…]
Finally, we have a large-scale structure that exhibits additional signatures. The emergent clustering density
$\rho_{\text{TD}}(k,a)=\Delta\mu(k,a)\bar{\rho}\delta(k,a)$ (96)
[…]
5 Law of Universal Arbitrage Equilibrium
We begin by recalling that the horizon entropy and temperature derived teleodynamically without assuming the Gibbons-Hawking form a priori are given by
$S_{\rm TD}(t)=\frac{C_{\rm TD}(t)}{H(t)^{2}}+\int^{t}\sigma_{\rm TD}(t^{\prime})\,dt^{\prime},\qquad T_{\rm TD}(t)=\left(\frac{\partial S}{\partial E_{H}}\right)^{-1}\simeq\frac{2\pi M_{P}^{2}}{C_{\rm TD}(t)}H(t)$ (97)
where the Misner-Sharp energy inside the Hubble sphere is $E_{H}=\rho_{\rm eff}V$, with $V=\tfrac{4\pi M_{P}^{2}}{H}$, the function $C_{\rm TD}(t)$ encodes the microstate density on the horizon and therefore plays the role of a cell entropy per correlation area, and $\sigma_{\rm TD}(t)\geq 0$ is the non-equilibrium entropy production generated by global teleodynamic memory. Taking a time derivative here would hence give us
$\dot{S}_{\rm TD}=-\frac{2C_{\rm TD}\dot{H}}{H^{3}}+\frac{\dot{C}_{\rm TD}}{H^{2}}+\sigma_{\rm TD}$ (98)
The energy flux across the apparent horizon produced by the effective bulk fluid $(\rho_{\rm eff},p_{\rm eff})$ follows the standard FRW prescription and is
$\dot{Q}=-\frac{4\pi}{H^{2}}(\rho_{\rm eff}+p_{\rm eff})$ (99)
Imposing a generalized Clausius relation with the teleodynamic temperature, where the usual horizon work term is absorbed by working with the apparent horizon energy flux, gives us
$\dot{Q}=T_{\rm TD}\,\dot{S}_{\rm TD}$ (100)
Substituting all ingredients into this balance law yields a teleodynamic form of the Raychaudhuri equation
$\dot{H}=-\frac{\rho_{\rm eff}+p_{\rm eff}}{2M_{P}^{2}}+\frac{H}{2}\frac{\dot{C}_{\rm TD}}{C_{\rm TD}}+\frac{H^{3}}{2C_{\rm TD}}\,\sigma_{\rm TD}$ (101)
where the numerical normalization of the first term has been chosen so that in the equilibrium limit with $C_{\rm TD}$ constant and $\sigma_{\rm TD}=0$ one recovers the standard general relativistic identity $2M_{P}^{2}\dot{H}=\rho_{\rm eff}+p_{\rm eff}$.
Now, if we go by multiplying the equation (101) by $2H$ and using $\tfrac{d}{dt}(H^{2})=2H\dot{H}$ together with the continuity equation $\dot{\rho}_{\rm eff}+3H(\rho_{\rm eff}+p_{\rm eff})=0$ then this leads us to
$\frac{d}{dt}(H^{2})=\frac{1}{3M_{P}^{2}}\dot{\rho}_{\rm eff}+H^{2}\frac{\dot{C}_{\rm TD}}{C_{\rm TD}}+H^{4}\frac{\sigma_{\rm TD}}{C_{\rm TD}}$ (102)
Integrating from a reference epoch $t_{\star}$ results in the teleodynamic Friedmann constraint as
$H^{2}(t)=\frac{\rho_{\rm eff}(t)}{3M_{P}^{2}}+\left[H^{2}(t_{\star})-\frac{\rho_{\rm eff}(t_{\star})}{3M_{P}^{2}}\right]+\int_{t_{\star}}^{t}\left(H^{2}\frac{\dot{C}_{\rm TD}}{C_{\rm TD}}+H^{4}\frac{\sigma_{\rm TD}}{C_{\rm TD}}\right)dt^{\prime}$ (103)
Note that this reduces to the standard Friedmann equation if the teleodynamic effects were negligible at $t_{\star}$ and collapses identically to $H^{2}=\rho_{\rm eff}/(3M_{P}^{2})$ when $C_{\rm TD}$ is constant and $\sigma_{\rm TD}$ vanishes.
The bias-functional description employed independently writes the effective energy density and pressure as
$\rho_{\rm eff}=\rho_{m}+\rho_{r}+\rho_{\rm TD},\qquad\rho_{\rm TD}=\alpha\overline{\Phi},\qquad p_{\rm TD}=\alpha\overline{\Pi}_{\rm TD}.$
In this language, equation (101) becomes
$\dot{H}=-\frac{\rho_{m}+\rho_{r}+p_{r}}{2M_{P}^{2}}-\frac{\rho_{\rm TD}+p_{\rm TD}}{2M_{P}^{2}}+\frac{H}{2}\frac{\dot{C}_{\rm TD}}{C_{\rm TD}}+\frac{H^{3}}{2C_{\rm TD}}\,\sigma_{\rm TD}$ (104)
revealing that two distinct teleodynamic mechanisms contribute to the expansion rate with a bulk component carried by $(\rho_{\rm TD},p_{\rm TD})=\alpha(\overline{\Phi},\overline{\Pi}_{\rm TD})$ and a horizon component encoded in $C_{\rm TD}(t)$ and $\sigma_{\rm TD}(t)$ representing the evolution of the correlation scale and the non-equilibrium memory production on the horizon. If late-time acceleration is primarily driven by bulk bias, then $C_{\rm TD}$ remains approximately constant and $\sigma_{\rm TD}$ approaches zero, with $w_{\rm TD}\approx-1$ emerging naturally when $\overline{\Pi}_{\rm TD}\simeq-\overline{\Phi}$. If instead part of the acceleration originates from the horizon sector, then the integral terms in the teleodynamic Friedmann relation are nonzero, and one infers small but finite values for $\dot{C}_{\rm TD}/C_{\rm TD}$ or $\sigma_{\rm TD}$ at low redshift. Since $\sigma_{\rm TD}\geq 0$, the last term in (101) always drives $H^{2}$ upward, temporarily mimicking an effective equation of state below 1-1 without introducing ghosts.
[…]
The bias-fluid form of the Friedmann equations is
$3M_{P}^{2}H^{2}=\rho_{m}+\rho_{r}+\rho_{\rm TD},\qquad-2M_{P}^{2}\dot{H}=\rho_{m}+p_{m}+\rho_{r}+p_{r}+\rho_{\rm TD}+p_{\rm TD},$
supplemented with the teleodynamic closure
$\rho_{\rm TD}=\alpha\overline{\Phi},\qquad p_{\rm TD}=\alpha\overline{\Pi}_{\rm TD}.$
From the first principles horizon-thermodynamic viewpoint, we have the non-equilibrium first law as
$\dot{E}_{H}=T_{\rm TD}\dot{S}_{\rm TD}-p_{\rm eff}\dot{V}_{H}$ (105)
with
$\dot{S}_{\rm TD}=-2C_{\rm TD}\frac{\dot{H}}{H^{3}}+\sigma_{\rm TD},\qquad\dot{E}_{H}=-4\pi M_{P}^{2}\frac{\dot{H}}{H^{2}},\qquad\dot{V}_{H}=-4\pi\frac{\dot{H}}{H^{4}}$ (106)
implies, after some algebra, the identity
$p_{\rm eff}=\frac{\kappa}{4\pi}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD},\qquad\kappa=\frac{2\pi M_{P}^{2}}{C_{\rm TD}}$ (107)
This expression exhibits two notable and important features. First, the purely geometric part associated with the area law cancels exactly, so that only the non-equilibrium entropy production $\sigma_{\rm TD}$ contributes to $p_{\rm eff}$. Second, in an expanding late-time Universe for which $\dot{H}<0$, a positive $\sigma_{\rm TD}$ automatically yields a negative pressure, producing a dark energy-like effect without introducing new fundamental fields.
Identifying $p_{\rm eff}$ with the teleodynamic pressure $p_{\rm TD}$ and invoking the bias closure gives
$\alpha\overline{\Pi}_{\rm TD}=\frac{\kappa}{4\pi}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD}=\frac{M_{P}^{2}}{2C_{\rm TD}}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD}$ (108)
and substituting this relation into the Raychaudhuri equation gives a direct expression for the teleodynamic energy density,
$\alpha\overline{\Phi}=-2M_{P}^{2}\dot{H}-(\rho_{m}+p_{m}+\rho_{r}+p_{r})-\frac{M_{P}^{2}}{2C_{\rm TD}}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD}$ (109)
These identities form the dictionary between the bias variables $(\overline{\Phi},\overline{\Pi}_{\rm TD})$ and the horizon-thermodynamic variables $(C_{\rm TD},\sigma_{\rm TD})$. The implications of this dictionary are immediate as well as with $\dot{H}<0$ in the late Universe and $\sigma_{\rm TD}\geq 0$, one obtains
$p_{\rm TD}=\alpha\overline{\Pi}_{\rm TD}=-\frac{M_{P}^{2}}{2C_{\rm TD}}\frac{H^{5}}{|\dot{H}|}\sigma_{\rm TD}<0$
showing that non-equilibrium teleodynamic entropy production necessarily gives us negative pressure and a combined expression for the EOS parameter follows
$w_{\rm TD}=\frac{p_{\rm TD}}{\rho_{\rm TD}}=\frac{-\dfrac{M_{P}^{2}}{2C_{\rm TD}}\dfrac{H^{5}}{\dot{H}}\sigma_{\rm TD}}{-2M_{P}^{2}\dot{H}-(\rho_{m}+p_{m}+\rho_{r}+p_{r})-\dfrac{M_{P}^{2}}{2C_{\rm TD}}\dfrac{H^{5}}{\dot{H}}\sigma_{\rm TD}}$ (110)
At late times where $p_{m}\approx 0$ and the radiation terms are negligible, the last term dominates the expression $w_{\rm TD}$ that approaches 1-1 as $\sigma_{\rm TD}$ decays while $\dot{H}\to 0$, reflecting the expected behavior of the attractor de Sitter. A minimal teleodynamic parameterization consistent with equilibrium recovery can be easily obtained by choosing a positive entropy production law that vanishes as equilibrium is approached, for instance
$\sigma_{\rm TD}(t)=2\xi\frac{-\dot{H}}{H^{3}},\qquad\xi>0$ (111)
With this choice, one finds
$p_{\rm TD}=-\frac{\kappa}{2\pi}\xi H^{2}=-\frac{M_{P}^{2}}{C_{\rm TD}}\xi H^{2},\qquad\rho_{\rm TD}=\frac{3M_{P}^{2}H^{2}}{8\pi}\Big(\cdots\Big)+\frac{M_{P}^{2}}{C_{\rm TD}}\xi H^{2}$ (112)
so both $p_{\rm TD}$ and $\rho_{\rm TD}$ scale as $H^{2}$ and the equation of state approaches 1-1 smoothly at late times. This produces a dynamically evolving DE-like component without new particles, consistent with the upward push in $H(z)$ required to ease the $H_{0}$ tension.
The emergent temperature $T_{\rm TD}=(2\pi M_{P}^{2}/C_{\rm TD})H$ shows that if cosmology selected $C_{\rm TD}=8\pi^{2}M_{P}^{2}$, the standard Gibbons-Hawking normalization would be recovered, and more generally, $C_{\rm TD}$ becomes an observable calibration of teleodynamic microphysics and is jointly constrained with the non-equilibrium parameter $\xi$ by background and perturbation data. Since $p_{\rm TD}$ is proportional to $\sigma_{\rm TD}$ and $\sigma_{\rm TD}$ is driven by nonlocal memory degrees of freedom, any measurement of $w_{\rm TD}(z)$ determines the degree of nonequilibrium production. This would therefore fix the strength of the force side correction in the Euler or Vlasov sector. This way, a single set of teleodynamic parameters can simultaneously explain late-time background uplift, which would help resolve the $H_{0}$ tension, and intermediate scale growth suppression, which would help address the $S_{8}$ tension.
Crucially, equating the bias fluid and horizon-thermodynamic constructions of teleodynamic cosmology demonstrates, in a mathematically explicit way, that global teleodynamic memory corresponds to the negative-pressure component usually identified as dark energy, while local teleodynamic memory produces clustering and behaves as dark matter. The teleodynamic pressure
$p_{\rm TD}=\frac{M_{P}^{2}}{2C_{\rm TD}}\frac{H^{5}}{\dot{H}}\sigma_{\rm TD}$
depends exclusively on horizon-level quantities such as $C_{\rm TD}$, $\sigma_{\rm TD}$, $H$, and $\dot{H}$ and none of the spatially varying components of the teleodynamic memory field enter this expression, indicating that cosmic acceleration arises only from the homogeneous part of the memory sector. Global memory, therefore, generates dark energy behavior while local memory enters through spatial gradients such as $\alpha\nabla\varphi$ and drives clustering, thereby reproducing dark matter like dynamics.
The form of the pressure above also shows that because $\sigma_{\rm TD}\geq 0$ and $\dot{H}<0$ in an expanding Universe, the negative pressure driving acceleration emerges if and only if the accessible microstates at the horizon increase. Teleodynamics, therefore, provides a thermodynamic demonstration and a law.
The Law of Universal Arbitrage Equilibrium: The late-time acceleration follows from the teleodynamic tendency of the Universe to evolve toward a state where the expansion no longer produces significant changes in the horizon-accessible microstate count. As the growth of these microstates stabilizes, the dynamics naturally approach an asymptotic accelerated phase, and dark energy is the macroscopic signature of this stabilization process.
6 Fates of a Teleodynamic Universe
The teleodynamic Friedmann system is first written in its bias–functional form, where the effective energy density and pressure are given by
$3M_{P}^{2}H^{2}=\rho_{m}+\rho_{r}+\rho_{\rm TD},\qquad-2M_{P}^{2}\dot{H}=\rho_{m}+p_{m}+\rho_{r}+p_{r}+\rho_{\rm TD}+p_{\rm TD},$
with the teleodynamic closure
$\rho_{\rm TD}=\alpha\,\overline{\Phi}(t),\qquad p_{\rm TD}=\alpha\,\overline{\Pi}_{\rm TD}(t).$
The various future scenarios of a teleodynamic universe follow directly from the behavior of $\overline{\Phi}(t)$ and from the corresponding evolution of $\rho_{\rm TD}(t)=\alpha\overline{\Phi}(t)$ and the equation of state parameter.
The first possibility arises when the homogeneous memory saturates as the structure formation approaches completion, as in that case $\overline{\Phi}(t)$ tends to a constant and the teleodynamic energy density approaches a positive constant value
$\rho_{\rm TD}(t)\to\rho_{\rm TD,\infty}=\alpha\,\overline{\Phi}_{\infty},\qquad w_{\rm TD}\to-1$ (113)
[…]
A second outcome occurs when homogeneous memory continues to grow slowly at late times, and this may happen if large-scale structures keep stretching, voids expand coherently, or long-range correlations persist, which ends up giving us
$\dot{\overline{\Phi}}(t)>0,\qquad w_{\rm TD}(t)<-1$ (114)
The teleodynamic energy density then grows with time,
$\rho_{\rm TD}(t)\propto\overline{\Phi}(t)$ (115)
producing a mild phantom-like expansion whose origin is statistical rather than exotic. Depending on how long $\overline{\Phi}(t)$ grows, the future evolution may approach a smoothly super-accelerating fate or a Big Rip-type divergence [99, 100, 101, 102], although without violating standard energy conditions.
A third possible regime emerges if the Universe eventually becomes too dilute for structures to maintain coherent tidal memory, as in that limit, both the homogeneous and spatial components of memory decay
$\overline{\Phi}(t)\to 0,\qquad\varphi(x,t)\to 0$ (116)
so that the teleodynamic contribution vanishes
$\rho_{\rm TD}(t)\to 0,\qquad w_{\rm TD}\to 0$ (117)
With the acceleration term disappearing, the Universe transitions into a coasting type phase. This is one where the asymptotic expansion behaves as $a(t)\sim t$, a huge asymptotic scale factor that is not infinite, which would be a Long Freeze scenario [103, 104, 105, 106]. Teleodynamic memory requires structure to exist, but if structure evaporates itself, then so does the memory that drives acceleration.
A fourth fate becomes possible if the evolution of memory reverses sign at late times. If large-scale correlations decay in such a way that the homogeneous bias becomes negative, one has
$\overline{\Phi}(t)<0,\qquad\rho_{\rm TD}(t)=\alpha\,\overline{\Phi}(t)<0$ (118)
This produces an additional attractive contribution to the cosmic dynamics. This is interesting as it can potentially trigger re-collapse [107, 108, 109, 110] even in geometrically flat space, as such a scenario requires large-scale anti-correlation that decreases horizon-level entropy carried by structure. This is unlikely in comparison to the other scenarios but is permitted within the teleodynamic framework.
A final possibility is a teleodynamic cyclic evolution, which is quite interesting, too. If memory increases during phases of structure buildup but decreases during ultra late homogenization, then
$\overline{\Phi}(t):\quad\text{growth}\;\longrightarrow\;\text{decay}$ (119)
This would lead either to an accelerated expansion or to a period of near-constant acceleration, followed by an eventual deceleration as memory fades. The universe may then contract, and the sequence can repeat, and such cycles [111, 112, 113] require no scalar fields, no ekpyrotic potentials, no ghosts, and no additional ingredients. The dynamics here arises purely from the time dependence of the teleodynamic memory kernel. The main point here is that teleodynamics allows for a wide variety of final fates for the universe.
In passing, we should also now mention what the possible fates for the universe are for the bias functional in the form given in (60). The first natural fate for this ansatz is a de Sitter–like future, since although the functional contains a term that grows as $a^{n}$ at late times, the physical meaning of the parameter $\epsilon$ is the ability of the structure to continue generating new memory. In this regime, the homogeneous bias asymptotes to a constant, $\bar{\Phi}(t)\to\bar{\Phi}_{\infty}$, and the teleodynamic energy density settles to a fixed positive value. The teleodynamic equation of state then tends toward $w_{\rm TD}\to-1$, driving the Universe into a de Sitter-like attractor. The second class of asymptotic behaviors arises when the teleodynamic memory continues to build indefinitely, corresponding to a persistent, non-vanishing $\epsilon$ even at ultra late times. In this case, the term $a^{n}$ grows monotonically without bound, causing the homogeneous bias to increase without saturation. If the rate of memory accumulation is sufficiently strong, the energy density diverges in finite cosmic time, resulting in a Big Rip-type future in which the scale factor, Hubble rate, and teleodynamic bias all blow up.
7 What is Memory in the Teleological Perspective?
It is helpful to discuss the subtleties of what really constitutes “memory” in the teleodynamic perspective, given all the formulations we have discussed so far.
In the teleodynamic framework, memory is not a psychological or informational construct but a physical property of an evolving cosmic system. It represents the persistent imprint left on the large-scale gravitational field by the history of structure formation, matter flows, and tidal interactions across cosmic timescales. Because gravity in an expanding universe is long-range and non-ergodic, the distribution of matter does not “forget” its past configurations, which is to say that we cannot treat the physics here to be completely Markovian. Instead, every merger, filamentary collapse, void expansion, or tidal alignment leaves traces in the statistical and systemic organization of the cosmic web, and these persistent features accumulate into a global bias functional, which we here refer to as cosmological memory.
From a dynamical perspective, memory encodes how the Universe departs from the assumptions of perfect equilibrium that underlie standard statistical mechanics. This is exemplified by the fact that a collisionless, self-gravitating system never fully thermalizes, and its phase-space distribution maintains long-lived correlations that cannot be erased by expansion alone. These correlations are precisely what teleodynamics captures. As the Universe evolves, preferred organizational patterns survive Hubble timescales and feed back into both the background expansion and the growth of structure. In this sense, we say that memory arises as an emergent, non-local statistical field that quantifies how past gravitational organization biases present evolution.
Cosmologically speaking, this memory becomes especially important once structure formation becomes nonlinear, which is the case once structures have been forming for an extended period of time, as in the current age of the universe. Filaments, clusters, and voids act as reservoirs of persistent tidal information, storing the integrated effects of billions of years of gravitational interactions. This accumulated organization cannot be described by local density alone; instead, it manifests as a global functional of the cosmic web. Viewed this way, cosmological memory is not an additional field or exotic substance introduced by hand, but is instead the inevitable statistical and systemic outcome of a Universe that has evolved under long-range gravity, formed a hierarchy of structures, and retained correlations across cosmic timescales. Teleodynamics formalizes this intuition into a predictive framework as the accumulated memory of structure formation influences large-scale dynamics, shapes late-time acceleration, and modulates the growth of perturbations. In this sense, memory is the physical bridge connecting the microscopic history of gravitational organization to the macroscopic observables that define modern cosmology.
8 Summary
We present a novel, potential game-theory-inspired statistical framework for cosmology, called Cosmological Teleodynamics. Our approach applies a maximum-caliber principle to cosmic histories, introducing a teleodynamic bias functional, $\Phi$, that encodes structural and evolutionary memory, yielding an additional drift force term $\alpha\nabla\Phi$ in the Boltzmann and Euler equations. In this framework, the dark sector, cosmic acceleration, large-scale structure formation, and the resolution of observational tensions emerge naturally from nonlocal memory effects and persistent self-organization in the gravitating Universe, rather than from new fundamental particles or fields.
The bias functional decomposes into $\Phi=\bar{\Phi}+\varphi$, with $\bar{\Phi}$ generating an effective homogeneous energy density $\rho_{\text{TD}}$ that naturally drives late-time acceleration, thus explaining dark energy. The inhomogeneous component $\varphi$ modifies the Poisson equation through an emergent clustering density $\rho_{\text{TD}}(k,a)$, which reproduces dark matter-like behavior, including rotation curves, lensing, cluster dynamics, and scale-dependent growth. Different response kernels $K(k,a)$ allow teleodynamic “dark matter” to mimic CDM, WDM, SIDM, etc., as well as generate anisotropic, environment-dependent halo signatures aligned with the cosmic web. This framework can help alleviate current observational tensions, particularly the $H_{0}$ and $S_{8}$ discrepancies.
It also predicts novel phenomena that cannot arise from conventional particle dark matter or scalar-field dark energy, including anisotropic velocity fields and environment-dependent halo signatures. Teleodynamic entropy emerges from microstate counting at the Hubble horizon, revealing a fundamentally nonequilibrium origin of cosmic entropy without invoking black hole thermodynamics. The current expansion epoch of the universe can be interpreted as it evolving towards an arbitrage equilibrium, in which it tries to maximize the global effective utility of accessible microstates in every region, thereby giving rise to the Law of Universal Arbitrage Equilibrium. The teleodynamic setup of cosmology can also support different fates for the universe, as different forms of the bias functional can lead to distinct long-term futures.
Cosmological Teleodynamics thus provides a unified, emergent, and observationally testable alternative to the standard dark sector paradigm. Rather than invoking new particles, it roots cosmological phenomena in the intrinsic statistical and systemic structure of cosmic memory, revealing that the Universe itself operates as a giant potential game.
Acknowledgments
The work of OT is supported in part by the Vanderbilt Discovery Doctoral Fellowship. The authors thank Robert Scherrer and Abraham Loeb for their very helpful comments on the work. VV thanks Julia Velkovska and her colleagues in the Department of Physics and Astronomy at Vanderbilt University for their kind invitation to present the 2025 Wendell G. Holladay Lecture, which resulted in this paper.